OpenStudy (anonymous):

limit as x approaches zero of quantity negative six plus x divided by x to the fourth power.

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

\(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~0}\frac{6+x}{x^4}}\) this?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

actually negative 6

OpenStudy (anonymous):

\(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~0}\frac{6+x}{x^4}}\) \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~0}\frac{6}{x^4}+\lim_{x \rightarrow ~0}\frac{x}{x^4}}\) so I don't think you will get anything defined out of the limit

OpenStudy (anonymous):

wait, approaches -6?

OpenStudy (anonymous):

oh, -6+x! Doesn't matter

OpenStudy (anonymous):

you would be then getting \(\large \displaystyle \lim_{x \rightarrow ~0}\frac{-6}{x^4}+\frac{x}{x^4}\) and still DNE

OpenStudy (anonymous):

oh okay thanks!

OpenStudy (anonymous):

Wait so which one is it?

OpenStudy (anonymous):

it is infinity not zero, i think

Vocaloid (vocaloid):

I don't believe L'Hopital's rule applies, -6/0 doesn't count as an indeterminate form

OpenStudy (anonymous):

no she meant part 2, but there x's xcancel

OpenStudy (anonymous):

What is that rule though?

OpenStudy (anonymous):

that is to differentiate top and bottom, IF you get 0/0 or ∞/∞, when you plug in the value that x approaches into the limit

OpenStudy (anonymous):

this is L'Hospital's Rule

OpenStudy (anonymous):

Could you guys help me with this?

OpenStudy (anonymous):

So for example, I have: \(\large \displaystyle \lim_{x \rightarrow ~0}\frac{\sin(x)}{x}\) and there you would apply this rule (can you tell me why?)

OpenStudy (anonymous):

i don't see a reason for one-sided limit not to exist (unless the function is totally not on that interval, or if it goes into infinity - asymptote)

OpenStudy (plasmataco):

Well x=2 is an asymptote

Vocaloid (vocaloid):

@phunish well, it's asking for the limit as the function approaches x = 2 from the left, any ideas?

OpenStudy (anonymous):

ok, you see that the graph has two parts (two sticks :D) right/

OpenStudy (anonymous):

?